OpenStudy (anonymous):

A jar holds 15 red pencils and 10 blue pencils. One pencil is chosen and the replaced. A second pencil is then chosen. What is the probability of drawing two red pencils from the jar?

OpenStudy (anonymous):

multiply the probabilities together. probability first one is red is \(\frac{15}{25}=\frac{3}{5}\) and probability second one is red is also \(\frac{3}{5}\) so your job is to compute \[\frac{3}{5}\times \frac{3}{5}\]

OpenStudy (anonymous):

So it does not change since it was replaced?

OpenStudy (anonymous):

it does not change because you have 15 red pencils and 25 total for both draws yes, because it was replaced

OpenStudy (anonymous):

Thank you!